OpenStudy (anonymous):

Find the slope of the line that passes through the given pair of points. (−a + 1, b − 3) and (a + 1, −b)

Directrix (directrix):

slope = [(b - 3) - (-b)] / [(-a +1) - (a + 1)] = ([b - 3 + b)] / [ -a +1 -a -1 ] = (2b - 3) / (-2a) = -[ (2b - 3]) /[ -(-2a)] =( -2b + 3)/2a = (3 - 2b) / (2a)